Calculate the terminal voltage for a battery with an internal resistance of 0.950 Ω and an emf of 7.00 V when the battery is connected in series with an 750-Ω resistor.
Answer #1

net current through circuit

i = E/ (R+r) = 7 / ( 0.95 + 750)

i = 0.0093215 A

terminal voltage

V = E - iR

V = 7 - 0.0093215* 0.95

V = 6.91144 V

========

Comment before rate in case any doubt, will reply for sure.. goodluck